Each of the ten volumes of the collected works of Theodore Sturgeon is available in paperback for $\$$15 or in hardcover for $\$$25. Theresa buys a copy of each of the ten volumes for a total of $\$$220. How many hardcover volumes did she buy?
Answer by Boreal(15235)

P=x
H=10-x
x*15+(10-x)25=220
15x+250-25x=220
-10x=-30
x=3 paperback@$15=$45
10-x=7 hardcover@$25=$175ANSWER
They add to $220.